OpenStudy (anonymous):

I have an unlimited supply of standard 6-sided dice. What's the fewest number of dice that I have to simultaneously roll to be at least 90% likely to roll at least one 6?

OpenStudy (mimi_x3):

hmmmmmm

OpenStudy (mimi_x3):

the site sucks today ... keeps kicking me out

OpenStudy (anonymous):

That's alright, take your time. I just want to make sure I understand this

OpenStudy (mimi_x3):

Ok so basically we are add all the probabilities of landing on a 6 until we reach 90% or .9 So the probability of landing on a 6 is 1/6 do you agree?

OpenStudy (anonymous):

Yes, the probability is 1/6 or 0.16666667

OpenStudy (mimi_x3):

So the probability of landing on 6 the first time is 1/6 The probability of landing on 6 a second is 1/6 the probability of landing on a 6 the third is 1/6 So basically is P(6)=1/6+1/6+1/6 .... = .9 So basically the formula can be arranged as follows x(1/6)=.9 Solve for x x represents how many times you roll the die Do you follow?

OpenStudy (anonymous):

Oh ok! I get it

OpenStudy (anonymous):

so we then divide 0.9 by 1/6? Then round up to the nearest whole number?

OpenStudy (mimi_x3):

yup :)

OpenStudy (anonymous):

Thanks so much!

OpenStudy (mimi_x3):

Do you get why we added all the probabilities?

OpenStudy (anonymous):

Well I understand we do not multiply them, because multiplying them would get us the probability of rolling a 6 each time

OpenStudy (anonymous):

So we add them together because we just want one or more 6

OpenStudy (mimi_x3):

yaaaaa :D

OpenStudy (mimi_x3):

Ya get the concept so we are good